OpenStudy (anonymous):

If Zack makes 1½ quarts of lemonade, then he uses 6 lemons. If Zack makes 1½ quarts of lemonade, then he makes 4 servings. Zack uses 5 lemons. Which is a valid conclusion?

OpenStudy (anonymous):

6/4 = 5/x (6 lemons to 4 servings = 5 lemons to x servings) cross multiply (6)(x) = (5)(4) 6x = 20 x = 20/6 x = 3 1/3 5 lemons makes 3 1/3 servings :) oh...wait...are you asking for the number of servings or how many quarts ?

OpenStudy (anonymous):

is it asking for # of servings or how many quarts are used with 5 lemons ?

OpenStudy (anonymous):

@whpalmer4 ....can you take a look at this

OpenStudy (whpalmer4):

I'm not seeing what I'm supposed to choose between as a valid conclusion...

OpenStudy (whpalmer4):

Is that an exact reproduction of the wording of the problem?

OpenStudy (anonymous):

I don't know...it is not my question....I am just gonna leave it like it is and if the asker wants the quarts per 5 lemons, then they will tag me

OpenStudy (anonymous):

thanks for looking